OpenStudy (anonymous):

(2x+11)/3(x+1)2=A/x+1+B/(x+1)2 find the values of constant A and B

OpenStudy (anonymous):

so we (2x+11)/3 =A(x+1) +B or 2/3x +11/3 = Ax+A+B thus equating coefficient of x we have A=2/3 and A+B=11/3

OpenStudy (anonymous):

so simply here we know A=2/3 plugging in for A=2/3 in A+B=11/3 we get B=3

OpenStudy (hba):

@Hero You want me to ?